OpenStudy (shubhamsrg):

p(x) is a polynomial with real coefficients such that for some positive real numbers c,d and for all natural numbers n, we have c |n|^3 <= |p(n)| <= d |n|^3 Prove p(x) has a real zero

OpenStudy (shubhamsrg):

I have the solution, its a little one, but i don;t understand that :|

mathslover (mathslover):

Please post the solution so that we can have a look too. :)

OpenStudy (shubhamsrg):

it simply says, " Obviously p(x) has degree 3, hence the result. " o.O how can we say it has degree 3 like that? :O

OpenStudy (anonymous):

If the degree is more than 3 then \[ \lim_{n\to \infty} \frac {|p(n)|}{n^3} =\infty \]

OpenStudy (anonymous):

If the degree is less than 3 then \lim_{n\to \infty} \frac {|p(n)|}{n^3} =0

OpenStudy (anonymous):

So to satisfy both inequalities, p has to have degree 3

OpenStudy (anonymous):

\[\lim_{n\to \infty} \frac {|p(n)|}{n^3} =0 \] If the degree of p is less than 3